OpenStudy (anonymous):

Easy probability question... A container holds 13 red marbles, 9 blue marbles, and 16 green marbles. You will randomly select two marbles without replacement. How many ways can you select the marbles? What is the probability that you select 2 red marbles? What is the probability that you select a green marble and then a red marble?

OpenStudy (anonymous):

for the first one im guessing its 2/38

OpenStudy (anonymous):

oh okay so what do i put for P?

OpenStudy (kropot72):

Sorry, my bad.

OpenStudy (anonymous):

lol so just forget about that equation?

OpenStudy (kenljw):

TM total marbles = R + B +G First draw P(R) = R/TM, P(B) = B/TM, P(G) = G/TM Then second draw you have to consider condition probabilities such as P(G|G) = (G - 1)/(TM-1) PG|B) = G/(TM-1) ect then use Baye's formual

OpenStudy (anonymous):

what does P stand for Probability?

OpenStudy (kenljw):

P(G) means the probability of green P(G|B) means the probability of green given previous draw bllue

OpenStudy (anonymous):

oh okay but there are a total of 38 marbles and for the first question is ask, How many ways can you select the marbles? it should be 2/38 cause you take 2 marbles out

OpenStudy (anonymous):

right?

OpenStudy (kenljw):

MT! is they were all different color For different colors MT!/(R!B!G!) It's divided by R! because it repeats that many times

OpenStudy (anonymous):

so what does MT! equal?

OpenStudy (primeralph):

why always marbles?

OpenStudy (kenljw):

MT to marbles = #R + #B +#G= R+B+G

OpenStudy (kenljw):

38! = 38X37X36 ... X1 very large and probably wont show up on calulator

OpenStudy (zarkon):

if you care about the order then there are 38*37 ways to pick 2 marbles

OpenStudy (kenljw):

Theres three colors

OpenStudy (zarkon):

there are 13*12 ways to pick two reds so the prob is (13*12)/(38*37)

OpenStudy (anonymous):

oh okay thanks both for taking the time to help me :)

OpenStudy (anonymous):

i got 151.89

OpenStudy (zarkon):

for what?

OpenStudy (anonymous):

(13*12)/(38*37)

OpenStudy (zarkon):

you might want to try that again

OpenStudy (kenljw):

Probabilities are always less than or equal to 1

OpenStudy (primeralph):

You can't get an answer greater than 1; at least not in this scope.

OpenStudy (zarkon):

parentheses are important

OpenStudy (anonymous):

so i have to put it in fraction

OpenStudy (reemii):

\(\frac{13\times 12}{38\times37}\) is what you have to compute.

OpenStudy (anonymous):

156/1406

OpenStudy (reemii):

that's is correct. (156/1406 = (13*12)/(38/37))

OpenStudy (anonymous):

oh okay thanks for the help :)

OpenStudy (anonymous):

umm who wants the medal?

OpenStudy (reemii):

oh i didn't work as much as the others.. :p thanks

OpenStudy (anonymous):

lol welcome, and i became of fan of the others who helped

OpenStudy (anonymous):

*a fan

OpenStudy (primeralph):